OpenStudy (anonymous):

how do u prove the identity of : tan x* cos^2 x -cot x* sin^ 2 x=0 ???

OpenStudy (usukidoll):

try using these identities \[\tan(x) = \frac{\sin(x)}{\cos(x)}, \cot (x) = \frac{\cos(x)}{\sin(x)}\]

OpenStudy (anonymous):

but wat do i do with sin^2 x

OpenStudy (usukidoll):

sin^2x or \[\sin^2(x) \] can also be written as \[\sin^2x = (sinx)(sinx) \] because that notation means to write sin x twice

OpenStudy (usukidoll):

Fair warning though \[\sin^2x \] and \[\sin2x \] ARE DIFFERENT! \[\sin^2x = (sin(x))(sin(x)) \] and \[\sin2x = 2sinxcosx \]

OpenStudy (usukidoll):

another hint: a cos x cancels for tan x* cos^2 x and a sinx cancels for cot x* sin^ 2 x afterwards the proof would be almost complete

OpenStudy (unklerhaukus):

after you substitute for tan and cot, terms will cancel.